# Enhancing Logical Expressiveness in Graph Neural Networks via Path-Neighbor Aggregation

### v2

@@ B.4 Baselines
The synthetic dataset is to verify the logical expressive power of the KG reasoning method based on GNN. We use two major categories of methods based on R-GNN and C-GNN as comparison baselines. The R-GNN-based methods include R-GCN (schlichtkrull2018modeling) and CompGCN (vashishth2019composition). The C-GNN-based methods include NBFNet (zhu2021neural) and RED-GNN (zhang2022knowledge). At the same time, it is also compared with EL-GNN (understanding) based on the labeling strategy.
@@
  • Methods based on representation learning:TransE (bordes2013translating), DistMult (yang2014embedding), RotatE (sun2019rotate), HAKE (zhang2020learning), RotH (chami2020low), ComplEx+RP (trouillon2016complex) and ConE (bai2021modeling).
  • Path-based methods: NeuralLP (yang2017differentiable), DRUM (sadeghian2019drum).
  • GNN-based methods: R-GCN (schlichtkrull2018modeling), CompGCN (vashishth2019composition), GraIL (teru2020inductive), RED-GNN (zhang2022knowledge), A*Net (zhu2024net), NBFNet (zhu2021neural), EL-GNN (understanding).
In inductive setting experiments, since representation-based methods need to learn specific entity representations, they cannot cope with the inductive setting. Therefore, path-based methods NeuralLP (yang2017differentiable), RuleN (meilicke2018fine), DRUM (sadeghian2019drum); and GNN-based methods NBFNett (zhu2021neural), RED-GNN (zhang2022knowledge), GraIL (teru2020inductive), CoMPILE (mai2021communicative) are used as comparison baselines. Since EL-GNN needs to label entities with entity representations, it cannot cope with the inductive setting.

### v3

@@ B.4 Baselines
The synthetic dataset is to verify the logical expressive power of the KG reasoning method based on GNN. We use two major categories of methods based on R-GNN and C-GNN as comparison baselines. The R-GNN-based methods include R-GCN (Schlichtkrull et al. 2018) and CompGCN (Vashishth et al. 2020). The C-GNN-based methods include NBFNet (Zhu et al. 2021) and RED-GNN (Zhang and Yao 2022). At the same time, it is also compared with EL-GNN (Qiu et al. 2024) based on the labeling strategy.
@@
  • Methods based on representation learning:TransE (Bordes et al. 2013), DistMult (Yang et al. 2015), RotatE (Sun et al. 2019), HAKE (Zhang et al. 2020), RotH (Chami et al. 2020), ComplEx+RP (Trouillon et al. 2016) and ConE (Bai et al. 2021).
  • Path-based methods: NeuralLP (Yang et al. 2017), DRUM (Sadeghian et al. 2019).
  • GNN-based methods: R-GCN (Schlichtkrull et al. 2018), CompGCN (Vashishth et al. 2020), GraIL (Teru et al. 2020), RED-GNN (Zhang and Yao 2022), A*Net (Zhu et al. 2023), NBFNet (Zhu et al. 2021), EL-GNN (Qiu et al. 2024).
In inductive setting experiments, since representation-based methods need to learn specific entity representations, they cannot cope with the inductive setting. Therefore, path-based methods NeuralLP (Yang et al. 2017), RuleN (Meilicke et al. 2018), DRUM (Sadeghian et al. 2019); and GNN-based methods NBFNett (Zhu et al. 2021), RED-GNN (Zhang and Yao 2022), GraIL (Teru et al. 2020), CoMPILE (Mai et al. 2021) are used as comparison baselines. Since EL-GNN needs to label entities with entity representations, it cannot cope with the inductive setting.